# Semiparametric KSD test: unifying score and distance-based approaches for goodness-of-fit testing

### v1

@@ 3.4 Bootstrap test: agnostic consistency guarantee
1 Obtain estimate $\hat{\theta}_{n}=\mathcal{E}(X)$;
2 Compute $T_{n}\equiv T_{n}(X,\hat{\theta}_{n})$ as in (SKSD);
3 for $b=1,2,\dots,B$ do
4   Generate bootstrap sample $\tilde{X}^{(b)}=(\tilde{X}_{1}^{(b)},\ldots,\tilde{X}_{n}^{(b)})\overset{\text{i.i.d.}}{\sim}\mathcal{L}_{\hat{\theta}_{n}}$;
5   Obtain bootstrap estimate $\tilde{\theta}_{n}^{(b)}=\mathcal{E}(\tilde{X}^{(b)})$;
6   Compute the resampled statistic $\widetilde{T}_{n}^{(b)}\equiv T_{n}(\tilde{X}^{(b)},\tilde{\theta}_{n}^{(b)})$.
7 end for
@@ 5   Obtain bootstrap estimate $\tilde{\theta}_{n}^{(b)}=\mathcal{E}(\tilde{X}^{(b)})$;
We want to point out that using parametric bootstrap is certainly not new in the literature. It has been a standard approach for goodness-of-fit testing [freedman1981bootstrapping, genest2008validity]. In the context of SKSD test, in addition to resolving the intractable distribution, there are two other critical reasons to favor the parametric bootstrap procedure. First, in the general context of GoF testing, it has been pointed out that other forms of bootstrap methods can fail including nonparametric bootstrap [bollen1992bootstrapping] and wild bootstrap [key2025composite, brueck2025composite]. Second, the proposed bootstrap procedure (Algorithm 1) is agnostic to the estimation procedure $\mathcal{E}$ and there is no need to estimate any nuisance parameter involved. We make the following remark on the computation complexity of the proposed bootstrap procedure.
@@
The computational complexity of Algorithm 1 is driven by three components: (1) the cost of the estimation procedure $\mathcal{E}$, denoted by $E(n)$ per fit with $n$ samples; (2) the cost of computing the test statistics $T_{n}$ and $\widetilde{T}_{n}$, which is $O(n^{2})$ per evaluation; and (3) the cost of generating bootstrap samples, denoted by $S(n)$ to obtain $n$ samples. Consequently, the total complexity is on the order of $O(B(E(n)+n^{2}+S(n)))$. In regimes where estimation or resampling is substantially more expensive (e.g., $E(n)\gg n^{2}$ or $S(n)\gg n^{2}$), the total cost can greatly exceed $O(Bn^{2})$. One possible remedy that avoids repeated resampling and refitting is to construct a Neyman-orthogonalized kernel [chernozhukov2022locally] so that the influence function no longer appears in (9) and wild bootstrap can be applied to achieve the validity [escanciano2024gaussian]. However, the resulting procedure has complexity $O(n^{3}+Bn^{2})$ (see Appendix C for details). Hence, there is a trade-off between the cost of computing the test statistic and the costs associated with estimation and resampling.
@@
Theorem 3 is a strong result in the context of bootstrap consistency. Unlike most of the existing bootstrap consistency results in kernel-based testing literature, Theorem 3 provides a universal weak limit regardless of how the data is generated, i.e., from the null or alternative hypothesis, despite the complexity of the asymptotic null distribution. The closest result to Theorem 3 is Theorem 5 in [key2025composite]. However, they only provide bootstrap validity under the null hypothesis when the MMD is used as the test statistic at the presence of a nuisance parameter. Therefore, the results cannot be used to analyze the power of the test. In contrast, our result allows us to understand the asymptotic behavior of the proposed test under both null and alternative hypotheses, providing particularly strong theoretical support for power analysis.
@@
To prove Theorem 3, we regard the result as a conditional version of weak convergence result in Theorem 2. The extra critical techniques used here are results on stable convergence [hausler2015stable], which provides a precise depiction of the weak convergence of the conditional distribution in the sense of Markov kernel convergence. We then show the proof strategy in Theorem 2 can be adapted to the conditional case, and be proved to hold stably. The property of martingale stable convergence leads to the target result.
@@
Corollary 1 guarantees the asymptotic validity of the SKSD test under a generic nuisance estimation procedure. In practice, when applying the proposed method with a particular class of estimators, analysts can readily verify Assumption 1, along with other standard regularity conditions.
@@ 4 Power analysis of SKSD test
With the asymptotic distribution of $T_{n}(X,\hat{\theta}_{n})$ and the bootstrap consistency result, we can now establish the asymptotic validity and consistency of the proposed bootstrap test $\phi_{n,\alpha}$.
Theorem 4 (Universal consistency against fixed alternative).
Suppose Assumptions 1 and regularity conditions in Appendix D hold. If $\inf_{\theta\in\Theta}\mathrm{KSD}(\mathcal{L}_{\theta},\mathcal{L})>0$ under $H_{1}$, then we have $\lim_{n\rightarrow\infty}\mathbb{P}_{H_{1}}[\phi_{n,\alpha}=1]=1$.
Theorem 4 shows that the proposed bootstrap test is consistent against any fixed alternative distribution that is separated from the null model in terms of KSD. This result is quite general as it holds for any UALE estimator $\hat{\theta}_{n}$ and does not require any further specification on the alternative distribution. We emphasize that the agnostic bootstrap consistency result under general hypotheses (Theorem 3) is what makes such power analyses possible. In practice, what is more interesting is the power behavior of the proposed test under local alternative models, which we will study in the following section. Throughout the rest of the section, we allow the law generating data, $\mathcal{L}$, to depend on the sample size $n$ and adopt the triangular array setup. We will use $\mathcal{L}_{n}$ to emphasize the dependence. Consider the following assumption.
@@
Suppoe either of the two conditions holds:
  1. a
    (Multiplicative local alternative, [jankova2020goodness]) Consider $H_{1,n}:\mathcal{L}_{n}\propto p_{\theta_{0}}(x)(1+h(x)/\sqrt{n})$ for some function $h(x):\mathbb{R}^{d}\to\mathbb{R}$. Further, suppose $h(X)$ has finite second moment under the null hypothesis, i.e., $\mathbb{E}_{\theta_{0}}[h^{2}(X)]<\infty.$
  2. b
    (Additive local alternative, [huber1965robust, niu2022distribution, chatterjee2025boosting]) Consider $H_{1,n}:\mathcal{L}_{n}=(1-\gamma/\sqrt{n})p_{\theta_{0}}(x)+(\gamma/\sqrt{n})g(x)$ for some density function $g(x):\mathbb{R}^{d}\to\mathbb{R}$ and constant $\gamma>0$. Further, suppose $\mathbb{E}_{\theta_{0}}\big[\big(\frac{g(X)}{p_{\theta_{0}}(X)}\big)^{2}\big]<\infty$ and denote $\mathcal{L}_{g}$ as the law with density $g$.
@@
The random variable $W_{1}\overset{d}{=}\sum_{u=1}^{\infty}\lambda_{u}(Z_{u}+\mu_{u})^{2}$ and $w_{1-\alpha}$ is the $(1-\alpha)$-quantile of $W$, where $\lambda_{u}$’s, $Z_{u}$’s and $W$ are defined in Theorem 2. If we write the eigenfunction corresponding to eigenvalue $\lambda_{u}$ of the operator $T_{\sigma}$ (defined in Theorem 2) as $\phi_{u}$, then we have $\mu_{u}=\mathbb{E}_{\theta_{0}}[S^{*}(\phi_{u},X,\theta_{0})h(X)]$ under Assumption 2a and $\mu_{u}=\gamma\mathbb{E}_{\mathcal{L}_{g}}[S^{*}(\phi_{u},X,\theta_{0})]$ under Assumption 2b. In particular, when there exists $u\in\mathbb{N}_{+}$ such that $\mu_{u}\neq 0$, we have $\beta>\alpha$.
Theorem 5 characterizes the asymptotic power function of our test under both multiplicative and additive local alternatives. The asymptotic power is fully characterized by the mean shift parameters $\mu_{u}$ and SKSD test $\phi_{n,\alpha}$ will have non-trivial power for local alternatives scaling with $1/\sqrt{n}$ as long as some $\mu_{u}\neq 0$. For multiplicative local alternatives, equivalently, this means the existence of some $f\in\mathcal{H}^{d}$ such that $\mathbb{E}_{\theta_{0}}[S^{*}(f,X,\theta_{0})h(X)]\neq 0$. This requires that under measure $\mathbb{P}_{\theta_{0}}$, tilting function $h(X)$ is correlated with at least one functional in the set $\{S^{*}(f,X,\theta_{0}):f\in\mathcal{H}^{d}\}$. For additive local alternatives, correspondingly, this is equivalent to the existence of some $f\in\mathcal{H}^{d}$ such that $\mathbb{E}_{\mathcal{L}_{g}}[S^{*}(f,X,\theta_{0})]\neq 0$. Overall, Theorem 5 shows the Pitman efficiency of our test. We demonstrate the finite-sample power of the SKSD test under these alternatives in Section K.
5 Application with SKSD test
In this section, we demonstrate the practical utility of our proposed SKSD test through three representative applications: testing normality in Section 5.1, determining order of kernel exponential family in Section 5.2, and detecting graphical structure in conditional Gaussian models in Section 5.3.
@@ 5.1 Testing normality
In this experiment, we consider four different data generating processes (DGPs) with Gaussian models the null model: (1) Gaussian distribution with different means $\mu$; (2) non-centered Student-$t$ distribution with varying degrees of freedom $\nu$; (3) mixture of two Gaussians with the same mixture weights and different mean shift $\delta$; and (4) non-centered generalized $\chi^{2}$ distribution with varying power parameter $\alpha$. Specifically, the DGPs are defined as follows: We refer to Appendix J.1 for the explicit forms of the distributions.
@@
To provide a comprehensive evaluation, we carry out simulations to examine finite-sample performance under various nuisance estimation strategies. The estimation and testing procedures used in this experiment are as follows. For estimation, we use minimum-KSD estimator [barp2019minimum] and MLE are used to estimate the parameters $(\mu,\sigma)$ in the null model $N(\mu,\sigma^{2})$. For testing, we consider the SKSD test, likelihood-ratio test (LRT), specialized normality tests, the Anderson-Darling test [anderson1952asymptotic] and the Lilliefors test [lilliefors1967kolmogorov], which are widely used in practice. The LRT in this context conincides with the well-known Vuong’s test [vuong1989likelihood]. Since we do not find a general method to calibrate the test invovled the nuisance parameter estimation, we use parametric bootstrap to approximate the null distribution of all tests.
@@
The sample size is set to $n=100$. The kernel used in the SKSD test is the Gaussian kernel with bandwidth selected based on the median heuristic [gretton2012kernel]. The nominal level $\alpha$ is set to be $0.05$. We keep the kernel and level choice fixed across all following experiments in this section.
Figure 1: Power curves of SKSD goodness-of-fit tests for Gaussian models under various null and alternative distributions. The data generating processes across the subfigures: top left: Gaussian distribution; top right: non-centered Student-$t$ distribution; bottom left: mixture of Gaussians; bottom right: non-centered generalized $\chi^{2}$ distribution. The black dashed line indicates the test level.

### v2

@@ 3.4 Bootstrap test: agnostic consistency guarantee
Obtain estimate $\hat{\theta}_{n}=\mathcal{E}(X)$;
Compute $T_{n}\equiv T_{n}(X,\hat{\theta}_{n})$ as in (SKSD);
for $b=1,2,\dots,B$ do
   Generate bootstrap sample $\tilde{X}^{(b)}=(\tilde{X}_{1}^{(b)},\ldots,\tilde{X}_{n}^{(b)})\overset{\text{i.i.d.}}{\sim}\mathcal{L}_{\hat{\theta}_{n}}$;
   Obtain bootstrap estimate $\tilde{\theta}_{n}^{(b)}=\mathcal{E}(\tilde{X}^{(b)})$;
   Compute the resampled statistic $\widetilde{T}_{n}^{(b)}\equiv T_{n}(\tilde{X}^{(b)},\tilde{\theta}_{n}^{(b)})$.
end for
@@
We want to point out that using parametric bootstrap is certainly not new in the literature. It has been a standard approach for GoF testing [34, 37]. In addition to resolving the intractable distribution, there are two other critical reasons to favor the parametric bootstrap procedure. First, in the general context of GoF testing, it has been pointed out that other forms of bootstrap methods can fail including nonparametric bootstrap [15] and wild bootstrap [58, 18]. Second, the proposed bootstrap procedure (Algorithm 1) is agnostic to the estimation procedure $\mathcal{E}$ and there is no need to estimate any nuisance parameter involved. We now make the following remark on the computational complexity of the proposed bootstrap procedure.
@@
The computational complexity of Algorithm 1 is driven by three components: (1) the cost of the estimation procedure $\mathcal{E}$, denoted by $E(n)$ per fit with $n$ samples; (2) the cost of computing the test statistics $T_{n}$ and $\widetilde{T}_{n}$, which is $O(n^{2})$ per evaluation; and (3) the cost of generating bootstrap samples, denoted by $S(n)$ to obtain $n$ samples. Consequently, the total complexity is on the order of $O(B(E(n)+n^{2}+S(n)))$. In regimes where estimation or resampling is substantially more expensive (e.g., $E(n)\gg n^{2}$ or $S(n)\gg n^{2}$), the total cost can greatly exceed $O(Bn^{2})$. One possible remedy that avoids repeated resampling and refitting is to construct a Neyman-orthogonalized kernel [22] so that the influence function no longer appears in (9) and classical wild bootstrap can be applied to achieve the validity [28]. However, the resulting procedure has complexity $O(n^{3}+Bn^{2})$ (see Appendix C for details). Hence, there is a trade-off between the cost of computing the test statistic and the costs associated with estimation and resampling.
@@
Theorem 3 provides a universal weak limit regardless of how the data are generated, i.e., from the null or alternative hypothesis, despite the complexity of the asymptotic null distribution. The closest result to Theorem 3 is Theorem 5 in [58]. However, they only provide bootstrap validity under the null hypothesis when the MMD is used as the test statistic at the presence of a nuisance parameter. Therefore, the results cannot be used to analyze the power of the test. In contrast, our result allows us to understand the asymptotic behavior of the proposed test under both null and alternative hypotheses, providing particularly strong theoretical support for power analysis.
@@
To prove Theorem 3, we interpret it as a conditional analogue of the weak convergence result established in Theorem 2. The key additional ingredient is the theory of stable convergence [49], which provides a rigorous framework for characterizing weak convergence of conditional distributions via convergence of Markov kernels. Within this framework, we show that the proof strategy used for Theorem 2 can be adapted to the conditional setting and established in a stable sense. Finally, the desired result follows from the martingale stable convergence result as in Lemma 2.
@@
Corollary 1 establishes the asymptotic validity of the SKSD test under a generic nuisance estimation procedure. In practice, when applying the proposed method with a specific class of estimators, analysts can readily verify Assumption 1 together with other standard regularity conditions.
@@ 4 Power analysis of SKSD test
With the consistency result of $T_{n}(X,\hat{\theta}_{n})$ (Theorem 1) and the bootstrap consistency (Theorem 3), we can now establish the power of the proposed bootstrap test $\phi_{n,\alpha}$. We begin with the consistency against fixed alternative distributions.
Theorem 4 (Universal power against fixed alternative).
Suppose Assumption 1 and regularity conditions in Appendix D hold. If $\inf_{\theta\in\Theta}\mathrm{KSD}(\mathcal{L}_{\theta},\mathcal{L})>0$ under $H_{1}$, then we have $\lim_{n\rightarrow\infty}\mathbb{P}_{H_{1}}[\phi_{n,\alpha}=1]=1$.
Theorem 4 shows that the proposed bootstrap test is consistent against any fixed alternative distribution that is separated from the null model in terms of KSD. This result is quite general as it holds for any UALE estimator $\hat{\theta}_{n}$ and does not require any further specification on the alternative distribution. We emphasize that the agnostic bootstrap consistency result under general hypotheses (Theorem 3) is what makes such power analyses possible.
In practice, what is more interesting is the power of the proposed test under local alternative models, which we will study now. Throughout the rest of the section, we allow the law generating data, $\mathcal{L}$, to depend on the sample size $n$ and adopt the triangular array setup. We will use $\mathcal{L}_{n}$ to emphasize the dependence. Consider the following assumption.
@@
Suppose either of the two conditions holds:
  1. (a)
    (Multiplicative local alternative, [55]) Consider the alternative laws $H_{1,n}:\mathcal{L}_{n}\propto p_{\theta_{0}}(x)(1+h(x)/\sqrt{n})$ for some function $h(x):\mathbb{R}^{d}\to\mathbb{R}$. Further, suppose $h(X)$ has finite second moment under the null, i.e., $\mathbb{E}_{X\sim\mathcal{L}_{\theta_{0}}}[h^{2}(X)]<\infty.$
  2. (b)
    (Additive local alternative, [53, 70, 20]) Consider the alternative laws $H_{1,n}:\mathcal{L}_{n}=(1-h/\sqrt{n})p_{\theta_{0}}(x)+(h/\sqrt{n})g(x)$ for some density function $g(x):\mathbb{R}^{d}\to\mathbb{R}$ and constant $h>0$. Further, suppose $\mathbb{E}_{X\sim\mathcal{L}_{\theta_{0}}}\big[\big(\frac{g(X)}{p_{\theta_{0}}(X)}\big)^{2}\big]<\infty$ and denote $\mathcal{L}_{g}$ as the law with density $g$.
@@
The random variable $W_{1}\overset{d}{=}\sum_{u=1}^{\infty}\lambda_{u}(Z_{u}+\mu_{u})^{2}$ and $w_{1-\alpha}$ is the $(1-\alpha)$-quantile of $W$, where $\lambda_{u}$’s, $Z_{u}$’s and $W$ are defined in Theorem 2. If we write the eigenfunction corresponding to eigenvalue $\lambda_{u}$ of the operator $T_{\sigma}$ (defined in Theorem 2) as $\phi_{u}$, then we have $\mu_{u}=\mathbb{E}_{X\sim\mathcal{L}_{\theta_{0}}}[S^{*}(\phi_{u},X,\theta_{0})h(X)]$ under Assumption 2(a) and $\mu_{u}=h\mathbb{E}_{X\sim\mathcal{L}_{g}}[S^{*}(\phi_{u},X,\theta_{0})]$ under Assumption 2(b). In particular, when there exists $u\in\mathbb{N}_{+}$ such that $\mu_{u}\neq 0$, we have $\beta>\alpha$.
Theorem 5 characterizes the asymptotic power function of our test under both multiplicative and additive local alternatives. The asymptotic power is fully characterized by the mean shift parameters $\mu_{u}$ and SKSD test $\phi_{n,\alpha}$ will have non-trivial power for local alternatives scaling with $1/\sqrt{n}$ as long as some $\mu_{u}\neq 0$. We make the following remark further discussing Theorem 5.
Remark 7 (Implication of Theorem 5).
For multiplicative local alternatives, equivalently, this means the existence of some $f\in\mathcal{H}^{d}$ such that $\mathbb{E}_{X\sim\mathcal{L}_{\theta_{0}}}[S^{*}(f,X,\theta_{0})h(X)]\neq 0$. This requires that under measure $\mathbb{P}_{\theta_{0}}$, function $h(X)$ is correlated with at least one element in the set $\{S^{*}(f,X,\theta_{0}):f\in\mathcal{H}^{d}\}$. For additive local alternatives, correspondingly, this is equivalent to the existence of some $f\in\mathcal{H}^{d}$ such that $\mathbb{E}_{X\sim\mathcal{L}_{g}}[S^{*}(f,X,\theta_{0})]\neq 0$. Overall, Theorem 5 shows the Pitman efficiency of our test. We demonstrate the finite-sample power of the SKSD test under these alternatives in Appendix K.
5 Application with SKSD test framework
In this section, we demonstrate the practical utility of our proposed SKSD test framework through three representative applications: testing normality in Section 5.1, determining order of kernel exponential family in Section 5.2, and detecting graphical structure in conditional Gaussian models in Section 5.3.
@@ 5.1 Testing normality
We consider four different data generating processes (DGPs) with Gaussian distribution being the null model: (1) Gaussian distribution with different means $\mu$; (2) non-centered Student-$t$ distribution with varying degrees of freedom $\nu$; (3) mixture of two Gaussians with equal mixture weights and mean shift parameter $\delta$; and (4) non-centered generalized $\chi^{2}$ distribution with varying signal parameter $\alpha$. We refer to Appendix J.1 for the explicit forms of the distributions. Setup (1) is under null while the others are under alternative.
Figure 1: Power curves of SKSD tests for Gaussian models under null and alternative DGPs. Top left: Gaussian distribution. Top right: non-centered Student-$t$ distribution. Bottom left: mixture of Gaussians. Bottom right: non-centered generalized $\chi^{2}$ distribution. The black dashed line indicates the test level.
@@
For estimation, we use minimum-KSD estimator [10] and MLE are used to estimate the parameters $(\mu,\sigma)$ under the null model $N(\mu,\sigma^{2})$. For testing, we consider the SKSD test, likelihood-ratio test (LRT), specialized normality tests, the Anderson-Darling test [5] and the Lilliefors test [61], which are widely used in practice. In this context, the LRT coincides with the well-known Vuong test [94], under the assumption that the alternative distribution family is known. We use parametric bootstrap to approximate the null distribution of all tests.
@@
The sample size is set to $n=100$. The kernel used in the SKSD test is the Gaussian kernel with bandwidth selected based on the median heuristic [42]. The nominal level $\alpha$ is set to be $0.05$. We keep the kernel and level choices fixed across all following experiments in this section.
The results are summarized in Figure 1, where “min-KSD + KSD” and “MLE + KSD” curves correspond to our proposed test. First, when data are generated under the null hypothesis, the SKSD test and the specialized normality tests (Anderson-Darling and Lilliefors) achieve the nominal level without noticeable inflation. The LRT-based test with either the minimum-KSD or MLE estimator is substantially conservative. Second, under the three alternatives, the SKSD tests exhibit power comparable to that of the specialized normality tests. In some settings, such as the Gaussian mixture model, the SKSD tests even outperform both the Anderson-Darling and Lilliefors tests. The power of the LRT-based tests, however, are quite unstable across different setups. Third, the choice of estimation algorithm $\mathcal{E}$ on the nuisance parameter does lead to differences in the power of the SKSD tests although the power gap is not substantial.
5.2 Order detection of kernel exponential model
Let $q$ be a reference density on $\mathbb{R}^{d}$, and let $\kappa:\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}$ be a reproducing kernel associated with RKHS $\mathcal{H}_{\kappa}$. The kernel exponential family model [19] is indexed by a function $f\in\mathcal{H}_{\kappa}^{d}$ and is given by
$p_{f}(x)\propto q(x)\exp(\langle f,\kappa(x,\cdot)\mathbf{1}_{d}\rangle_{\mathcal{H}_{\kappa}^{d}}),\quad\text{where}\quad x\in\mathbb{R}^{d}$
and we consider $\kappa$ to be the standard Gaussian kernel and $q(x)$ to be the density of $N(0,9)$. In this case, $p_{f}$ is parameterized by the functional element $f\in\mathcal{H}_{\kappa}^{d}$, and the normalizing constant is typically intractable, involving evaluation of high-dimensional integral. This model offers a nonparametric and hence flexible modeling framework, which has been widely studied in the recent machine learning literature [88, 90].
A common strategy in this line of work is to approximate the unknown data-generating density using a finite-rank representation with sufficient complexity, i.e., for some $r\in\mathbb{N}^{+}$ and basis functions $\{\phi_{\ell}\}_{\ell=1}^{r}$ of $\mathcal{H}_{\kappa}^{d}$, approximate $f$ by $f_{r}=\sum_{\ell=1}^{r}\theta_{\ell}\phi_{\ell}(\cdot)$ for some coefficients $\theta_{\ell}\in\mathbb{R}$. This leads to a finite-dimensional parameterization of the model, which is more amenable to statistical inference and computationally efficient. Accordingly, in this experiment we focus on GoF testing for finite-rank approximations of the kernel exponential family.
Experiment setup.
Throughout this simulation, we consider the density of data are generated by a linear subspace of RKHS: $\mathcal{P}_{r}\coloneqq\{p_{f_{r}}:f_{r}=\sum_{\ell=1}^{r}\theta_{\ell}\phi_{\ell}(\cdot),\theta_{\ell}\in\mathbb{R},\ell\in[r]\}$. Two settings are applied: (1) The DGPs are chosen in $\mathcal{P}_{2}$ with $\theta_{1}=10$ and various $\theta_{2}$ to simulate different levels of signal strength, while the null model hypothesis is fixed as $\mathcal{P}_{1}$; (2) The DGPs are chosen as distributions in $\mathcal{P}_{5}$ with random chosen coefficients, while we consider different null models: $\mathcal{P}_{r}$ for $r=1,2,3,4$. Noting that $\mathcal{P}_{r_{1}}\subseteq\mathcal{P}_{r_{2}}$ for any $r_{1}<r_{2}$, we simulate a progressive model selection scenario here, where the null model becomes increasingly complex and approaches the true data-generating distribution as $r$ increases.
We generate the sample using metroplis-adjusted Langevin algorithm (MALA, [80]) and set sample size $n=200$. We use the minimum-KSD estimator to estimate the parameters in the density model, which has become a standard approach in the literature [66, 64]. We refer the reader to Appendix J.2 for the complete simulation setup.